OpenStudy (anonymous):

simplify 42a-18b --------- 6

mathslover (mathslover):

\[\huge{7a-3b}\]

Parth (parthkohli):

There's a way to do this. Take 6 as a common factor, then factor 42a - 18b. The thing inside the bracket would be the answer.

Parth (parthkohli):

Or, you could do term-by-term division.

mathslover (mathslover):

we can write \(\huge{\frac{a-b}{c}=\frac{a}{c}-\frac{b}{c}}\)

mathslover (mathslover):

hence we will write it as \(\huge\frac{42a}{6}-\frac{18b}{6}\)\(\huge{=}\)\(\huge{7a-3b}\)